OpenStudy (anonymous):

The region bounded by y = ex, y = e, and the y-axis is revolved around the y-axis. The volume of the resulting solid is:

OpenStudy (mimi_x3):

Is it \(y = e^{x}\)?

OpenStudy (anonymous):

yes :)

OpenStudy (mimi_x3):

revolved around the \(y\)-axis so.. we change it to make \(x\) the subject i think..it has been a while that i have dont these.. \[ y = e^x => x = ln(y)\]

OpenStudy (mimi_x3):

well do you know what to do next?

OpenStudy (anonymous):

i think that it's pi * integral of (e-ln(y))^2

OpenStudy (mimi_x3):

yes!

OpenStudy (mimi_x3):

\[V = \pi\int\limits \ln^{2}(y)dy \]

OpenStudy (anonymous):

and the limits are 0 to 2.718 am i correct?

OpenStudy (mimi_x3):

e to 0